# Thermodynamic theory of square skyrmion lattice in tetragonal frustrated antiferromagnets

### v1

@@ 4 Phase diagrams and application to GdRu2Si2
These exchange parameters, roughly speaking, determine the overall scales of the axes in the $T-H$ phase diagrams. Particular “filling” of the ordered phase region is determined by weak interaction, in our case, biaxial momentum-dependent anisotropy and biquadratic exchange. Without them, our mean-field approach yields a trivial phase diagram consisting of TC and paramagnetic phases. The anisotropy helps us to move TC to low temperatures, and biquadratic exchange further promotes the multiple-modulated structures. We note that for some particular parameter choices, the phase diagrams discussed below correspond well to those obtained numerically in Ref. Hayami and Kato (2023). This further justifies our analytical approach.
@@
These estimations align with the numerical calculations of Ref. Bouaziz et al. (2022). Anisotropy of the pure dipolar nature gives Utesov (2021)
@@
Figure 3: Phase diagrams for simplified models, i.e., for which we include only one of the crucial ingredients. (a) Without the biquadratic exchange, but including anisotropy due to magnetodipolar interaction, the 2Q phase is stable only in a tiny region. At lower $T$, it loses competition to the transverse conical spiral. (b) For nonzero biquadratic exchange with $K=0.05$ K but without anisotropy, TC is stable in a large region of moderate fields. (c) The same as (b), but $K=0.1$ K. Larger $K$ allows substituting TC with vortical phase 2SXY and makes the square skyrmion lattice stable in the intermediate field region.
We continue playing with the phase diagrams, trying to obtain something similar to the diagram shown in Fig. 1a of Ref. Khanh et al. (2022). We combine magnetodipolar anisotropy and biquadratic exchange to stabilize 2SXY and suppress TC. We also enhance the anisotropy parameter $\Lambda_{1}$ to have the temperature of the phase transition between 2S and 1Q+1S about $40$ K using the single-ion easy-plane anisotropy with $Z=0.1$ K (it makes $\Lambda_{1}=0.15$ K and $\Lambda_{2}=0.2$ K). We arrive at the phase diagram shown in Fig. 4(a) by taking biquadratic exchange constant $K=0.03$ K, and $J_{\mathbf{k}_{\pm}}=J_{\mathbf{0}}$ yielding $\Lambda_{\pm}=1.9$ K. Note that the latter value is reasonable according to numerical calculations of Ref. Bouaziz et al. (2022). Smaller $\Lambda_{\pm}=0.95$ allows expanding the skyrmion lattice stability domain, see Fig. 4(b). Noteworthy that further decreasing of $\Lambda_{\pm}$ and increasing of $K$ leads to the emergence of the topologically nontrivial 2Q phase predicted in Ref. Wang et al. (2021). This effect is illustrated in Fig. 4(c) for $\Lambda_{\pm}=0.5$ K and $K=0.1$ K.
@@
The phase diagram shown in Fig. 4(a) has important similarities with that of Ref. Khanh et al. (2022). However, there are also several crucial differences. First, we note that according to this study and the early paper Garnier et al. (1996), there should be a phase transition at $H\sim 1$ T between two ordered phases for $T>40$ K. Second, the phase diagram for $H\parallel[100]$ unambiguously shows that for $h=0$, one of the phases corresponds to the structure we dub 2SYZ. Third, the presaturation phase for $H\parallel[100]$ is SDW with a modulated component polarized along $\hat{z}$. This indicates that the $\hat{z}$ axis is the easy one for $\mathbf{k}_{x,y}$ Utesov and Syromyatnikov (2021). Importantly, as discussed above, it leads to essentially different phase diagrams.
In Fig. 5(a), we show the result for the dipolar anisotropy with an additional easy axis term $Z=0.2$ K, $\Lambda_{\pm}=0.95$ K, and zero biaxial anisotropy. For such parameters, the single-modulated helicoid 1Q emerges via a first-order phase transition from 1S, and it is stable for weak $h$. A small value of $K$ leads to a shrinkage of the single-modulated phases’ stability region and stabilization of 1Q+1S for low temperatures. The result for $K=0.03$ K is shown in Fig. 5(b); other parameters are the same as above. Finally, by taking $K=0$, $\Lambda_{\pm}=0.5$ K, and enhancing the anisotropic exchange ($\Lambda_{1}\approx-0.15$ K and $\Lambda_{2}\approx 0.4$ K), we arrive at the phase diagram 5(c) showing expanded SkL stability region even without the biquadratic exchange.
@@
Note that the phase diagram shown in Fig. 5(c) captures several important features observed for GdRu2Ge2 in Ref. Yoshimochi et al. (2024). Evidently, in our study, we miss phases with harmonics characterized by additional modulation vectors (±$(\pm k/2,\pm k/2,0)$. However, even in a simplified calculation, we see that we can discuss phases I, IV, V, and VII using the notations of Ref. Yoshimochi et al. (2024). Our 1Q+1S structure seems to lose a competition to the hybrid phases II and III. Moreover, in Ref. Yoshimochi et al. (2024), the model with $K=0$ and close competition between structures with $\mathbf{k}_{x,y}$ and (±$(\pm k/2,\pm k/2,0)$ was proposed. According to our findings, the latter is essential for the stabilization of the square SkL in the significant part of the phase diagram.
[…]
5 Discussion
@@
We believe that the way we obtained the phase diagram shown in Fig. 6(c) reflects the physics of GdRu2Si2. It requires frustrated exchange interaction $J_{\mathbf{q}}$ with four maxima at ±$\pm\mathbf{k}_{x,y}$, much smaller “diagonal” exchange $J_{\mathbf{k}_{\pm}}$, moderate anisotropic exchange with easy axis along $\hat{z}$ and compass in-play term to stabilize Bloch-type structures, and moderate biquadratic exchange $K$. The latter should be larger than the isotropic Landau coefficient, i.e., $K\gtrsim b=BT_{\textrm{N}}$.
@@
There are certain discrepancies between our result Fig. 6(c) and Fig. 1a of Ref. Khanh et al. (2022). Here, the thermal fluctuation can play an important role. Using the Ginzburg-Levanyuk criterion Landau and Lifshitz (2013), we obtain
@@
Similar arguments apply to the square skyrmion lattice 2Q. We can “push” it to lower $T$s by increasing the $\Lambda_{\pm}$ parameter. However, thermal fluctuations are also a plausible solution, and a narrow wedge of the 2Q SkL phase at $4030\lesssim T<$ 40 K in Fig. 6(c) can be hard to distinguish experimentally. Moreover, the topological charge of magnetic unit cells relies on the criterion $2(p+s_{\pm})>m$ derived in the continuum spin density approximation. In the lattice picture, if the difference $2(p+s_{\pm})-m\equiv\delta s\ll 1$, then for the majority of spins have $s^{z}_{i}>0$ and only a few have $s^{z}_{i}<0$. On the mean field level, it can be understood by considering the spin structure (12) with the correction (13). For $m>0$, the minima of $s^{z}_{i}$ correspond to $\mathbf{k}_{x,y}\mathbf{R}_{j}=\pi+2\pi l_{x,y}$. However, due to the incommensurate $\mathbf{k}_{x,y}$, the minimum varies among magnetic unit cells, and only part of them $\propto\delta s$ have at least one spin with $s^{z}_{i}<0$. Then, even without fluctuations, in the regions of the phase diagram close to the 2Q trivial phase, the average density of topological charge $\langle n_{\textrm{sk}}\rangle\ll 1$. In simple words, it means that “robust” skyrmions providing a tangible topological Hall effect probably do not exist in the high-$T$ part of the phase diagram, despite the corresponding phase being, in principle, stable.
6 Summary and conclusion
Using the high-temperature Landau expansion of the free energy, we analytically discuss noncollinear (multiple-)modulated phases of tetragonal frustrated antiferromagnets observed experimentally. We show that the subtle inter-phase competition can be discussed in the framework of a model that includes frustrated exchange coupling, anisotropic exchange, and/or magnetodipolar interaction, single-ion anisotropy, biquadratic exchange, and Zeeman energy. For the exchange coupling, we only assume that it has four equivalent maxima at ±$\pm\mathbf{k}_{x}=(\pm k,0,0)$ and ±$\pm\mathbf{k}_{y}=(0,\pm k,0)$ in the reciprocal space. The anisotropies are combined into a momentum-dependent biaxial term. Phase diagram topology essentially depends on the corresponding axes hierarchy.
@@ 6 Summary and conclusion
Easy axes along the high-symmetry direction allow observation of a single-modulated spin-density wave 1S and helicoid 1Q, see Figs. 5(a), (b), and (c). Notably, 1Q can be stable at low $T$ only for negligible biquadratic exchange. Similarly to the in-plane easy axes case, square SkL is stable at moderate fields. Experimentally, we believe that the case of out-of-plane easy axes sheds some light on the phase diagram of GdRu2Ge2 Yoshimochi et al. (2024). However, the observed phase diagram in that paper is much more complicated due to a close competition between two sorts of modulation vectors, and requires a separate study.
For enhanced biquadratic exchange, we observe stabilization of SDW superpositions 2SYZ and 2SZZ at high $T$, see Figs. 6(a), (b), and (c). Notably, it can be argued that 2SYZ should emerge in GdRu2Si2 phase diagram [see Refs. Garnier et al. (1996); Khanh et al. (2022)]. Then, by choosing reasonable parameters, we obtain the phase diagram shown in Fig. 6(c), which seems to reproduce all crucial observations for GdRu2Si2 Khanh et al. (2020, 2022).
To conclude, we developed a simple yet fruitful mean-field approach for the high-temperature part of the phase diagrams of tetragonal frustrated antiferromagnets. Interplay of complex magnetic anisotropies and biquadratic exchange leads to a variety of possible phase diagrams. By comparing our findings with experimental phase diagrams of relevant materials GdRu2Si2  and GdRu2Ge2, we can draw important conclusions about the corresponding microscopic parameters. In particular, for the former we predict the relative strength of the biquadratic exchange of the order of tens of percent of the corresponding Heisenberg exchange, whereas for the latter the biquadratic exchange should be negligible in comparison with the anisotropic terms, which parallels numerics of Ref. Yoshimochi et al. (2024).
CRediT authorship contribution statement
Oleg I. Utesov: Conceptualization, Investigation, Methodology, Visualization, Writing – original draft, Writing – review & editing. Danila P. Budylev: Investigation, Visualization.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Appendix A Free energies of spin structures
Here, we discuss the details of the analytical calculations and present the corresponding results. The proposed phases are mostly consist of spin-density waves (S) and screw helicoids (Q) with modulation vectors $\mathbf{k}_{x}$ and $\mathbf{k}_{y}$. This way, we can discuss all the relevant phases proposed and observed in numerical calculations Hayami and Motome (2021); Wang et al. (2021); Hayami (2022); Hayami and Kato (2023), as well as distinguished experimentally Khanh et al. (2020, 2022). Importantly, we consider only incommensurate spin structures, allowing us to freely choose the phases of the corresponding constituents (up to certain limitations; see discussion of the 2Q structure below). Magnetic field is oriented along $\hat{z}$ and its main effect is the homogeneous magnetization component $m$.
@@ A.2 Double-modulated superposition of spin-density waves (2S)
This spin structure can be referred to as a “magnetic bubble crystal”. It was addressed in detail in Ref. Hayami and Kato (2023). In this case,
@@ A.7 Alternative 2Q ordering (2Q)
The possibility of this spin structure emergence was predicted in Ref. Wang et al. (2021). In real space, it is given by
@@ Appendix B Topological properties of multiple-modulated spin structures
$\rho_{\textrm{2Q}}(\mathbf{r})=-\frac{2s^{2}\left[\sqrt{2}p\left(\cos{\mathbf{k}_{x}\mathbf{r}}+\cos{\mathbf{k}_{y}\mathbf{r}}\right)+2m\cos{\mathbf{k}_{x}\mathbf{r}}\cos{\mathbf{k}_{y}\mathbf{r}}-q\frac{6+2\cos{2\mathbf{k}_{x}\mathbf{r}}+2\cos{2\mathbf{k}_{y}\mathbf{r}}-\cos{2\left(\mathbf{k}_{x}+\mathbf{k}_{y}\right)\mathbf{r}}-\cos{2\left(\mathbf{k}_{x}-\mathbf{k}_{y}\right)\mathbf{r}}}{2\sqrt{2}}\right]}{\left[2s^{2}(\sin^{2}{\mathbf{k}_{x}\mathbf{r}}+\sin^{2}{\mathbf{k}_{y}\mathbf{r}})+\left\{2m+\sqrt{2}p(\cos{\mathbf{k}_{x}\mathbf{r}}+\cos{\mathbf{k}_{y}\mathbf{r}})-2\sqrt{2}q\cos{\mathbf{k}_{x}\mathbf{r}}\cos{\mathbf{k}_{y}\mathbf{r}}\right\}^{2}\right]^{3/2}}$ (65)
@@
We plot densities of topological charge for 1Q+1S, 2Q, and 2Q for certain reasonable parameters in Fig. 7 for a domain containing $2\times 2$ magnetic unit cells. 1Q+1S structure has zero net topological charge. However, there are peculiar stripes of topological charge due to meron-antimeron spin structure Khanh et al. (2022); Wood et al. (2023). Square skyrmion lattice 2Q has $n_{\textrm{sk}}=\pm 1$, whereas the alternative structure 2Q has $n_{\textrm{sk}}=\pm 2$ Hayami and Motome (2021); Wang et al. (2021).

### v2

@@ IV Phase diagrams and application to $\text{Gd}\text{Ru}_{2}\text{Si}_{2}$
These exchange parameters, roughly speaking, determine the overall scales of the axes in the THT-H phase diagrams. Particular “filling” of the ordered phase region is determined by weak interaction, in our case, biaxial momentum-dependent anisotropy and biquadratic exchange. Without them, our mean-field approach yields a trivial phase diagram consisting of TC and paramagnetic phases. The anisotropy helps us to move TC to low temperatures, and biquadratic exchange further promotes the multiple-modulated structures. We note that for some particular parameter choices, the phase diagrams discussed below correspond well to those obtained numerically in Ref. [40]. This further justifies our analytical approach.
@@
These estimations align with the numerical calculations of Ref. [33]. Anisotropy of the pure dipolar nature corresponds to [26]
@@
We continue playing with the phase diagrams, trying to obtain something similar to the diagram shown in Fig. 1a of Ref. [31]. We combine magnetodipolar anisotropy and biquadratic exchange to stabilize 2SXY and suppress TC. We also enhance the anisotropy parameter $\Lambda_{1}$ to have the temperature of the phase transition between 2S and 1Q+1S about $40$ K using the single-ion easy-plane anisotropy with $Z=0.1$ K (it makes $\Lambda_{1}=0.15$ K and $\Lambda_{2}=0.2$ K). We arrive at the phase diagram shown in Fig. 4(a) by taking biquadratic exchange constant $K=0.03$ K, and $J_{\mathbf{k}_{\pm}}=J_{\mathbf{0}}$ yielding $\Lambda_{\pm}=1.9$ K. Note that the latter value is reasonable according to numerical calculations of Ref. [33]. Smaller $\Lambda_{\pm}=0.95$ allows expanding the skyrmion lattice stability domain, see Fig. 4(b). Noteworthy that further decreasing of $\Lambda_{\pm}$ and increasing of $K$ leads to the emergence of the topologically nontrivial 2Q phase predicted in Ref. [27]. This effect is illustrated in Fig. 4(c) for $\Lambda_{\pm}=0.5$ K and $K=0.1$ K.
@@
The phase diagram shown in Fig. 4(a) has important similarities with that of Ref. [31]. However, there are also several crucial differences. First, we note that according to this study and the early paper [37], there should be a phase transition at $H\sim 1$ T between two ordered phases for $T>40$ K. Second, the phase diagram for $H\parallel[100]$ unambiguously shows that for $h=0$, one of the phases corresponds to the structure we dub 2SYZ. Third, the presaturation phase for $H\parallel[100]$ is SDW with a modulated component polarized along $\hat{z}$. This indicates that the $\hat{z}$ axis is the easy one for $\mathbf{k}_{x,y}$ [41]. Importantly, as discussed above, it leads to essentially different phase diagrams.
In Fig. 5(a), we show the result for the dipolar anisotropy with an additional easy-axis term $Z=0.2$ K, $\Lambda_{\pm}=0.95$ K, and zero biaxial anisotropy. For such parameters, the single-modulated helicoid 1Q emerges via a first-order phase transition from 1S, and it is stable for weak $h$. A small value of $K$ leads to a shrinkage of the single-modulated phases’ stability region and stabilization of 1Q+1S for low temperatures. The result for $K=0.03$ K is shown in Fig. 5(b); other parameters are the same as above. Finally, by taking $K=0$, $\Lambda_{\pm}=0.5$ K, and enhancing the anisotropic exchange ($\Lambda_{1}\approx-0.15$ K and $\Lambda_{2}\approx 0.4$ K), we arrive at the phase diagram 5(c) showing expanded SkL stability region even without the biquadratic exchange.
@@
Note that the phase diagram shown in Fig. 5(c) captures several important features observed for GdRu2Ge2 in Ref. [22]. Evidently, in our study, we miss phases with harmonics characterized by additional modulation vectors $(\pm k/2,\pm k/2,0)$. However, even in a simplified calculation, we see that we can discuss phases I, IV, V, and VII using the notations of Ref. [22]. Our 1Q+1S structure seems to lose a competition to the hybrid phases II and III. Moreover, in Ref. [22], the model with $K=0$ and close competition between structures with $\mathbf{k}_{x,y}$ and $(\pm k/2,\pm k/2,0)$ was proposed. According to our findings, the latter is essential for the stabilization of the square SkL in a significant part of the phase diagram.
Now it is evident that to obtain a phase diagram similar to that of GdRu2Si2, we need to increase $K$. As shown in Appendix A, if $K>b$ then 1Q loses competition to 2SYZ. This scenario is plausible for the description of the experimental observations. Note that there is another condition: if $K>3b$ then 1S loses competition to 2SZZ, which, in its turn, also competes with 2SYZ, shrinking its region of stability. Moreover, strong biquadratic exchange leads to stabilization of SkL even at large $\Lambda_{\pm}$. We illustrate this discussion in Fig. 6. In panel (a), we take $K=0.2$ K, $\Lambda_{\pm}=1.9$ K, and $Z=0.2$ K. In this case, high-$T$ low-$h$ phases are 1S and 2SYZ, and 2Q is stable in the moderate fields region. Taking a larger $K=0.3$ K, we substitute 1S with 2SZZ, see Fig. 6(b). Finally, to push 2Q to low temperatures and get rid of competition between 2SYZ and 2SZZ, we take $K=0.2$ K, $\Lambda_{\pm}=5$ K, $Z=0.25$ K and arrive at the phase diagram shown in Fig. 6(c).
Figure 6: Phase diagrams for the easy axes along the high-symmetry direction $\hat{z}$ and enhanced biquadratic exchange. (a) $K\in(b,3b)$ leads to stabilization of 2SYZ and 1S at high-$T$ and low-$h$ part of the phase diagram. (b) If $K>3b$, instead of 1S, the double SDW structure 2SZZ emerges. (c) For $K=0.2$ K, weakened competition among $\mathbf{k}_{x,y}$ and $\mathbf{k}_{\pm}$ (increased $\Lambda_{\pm}$), the well-pronounced skyrmion phase can be stabilized at lower temperatures. It is argued that this type of phase diagram corresponds to the real one of GdRu2Si2.
V Discussion
@@
We believe that the way we obtained the phase diagram shown in Fig. 6(c) reflects the physics of GdRu2Si2. It requires frustrated exchange interaction $J_{\mathbf{q}}$ with four maxima at $\pm\mathbf{k}_{x,y}$, much smaller “diagonal” exchange $J_{\mathbf{k}_{\pm}}$, moderate anisotropic exchange with easy axis along $\hat{z}$ and compass in-plane term (in the form of, e.g., magnetodipolar interaction) to stabilize Bloch-type structures, and moderate biquadratic exchange $K$. The latter should be larger than the isotropic Landau coefficient, i.e., $K\gtrsim b=BT_{\textrm{N}}\sim 0.1$ K.
@@
There are certain discrepancies between our result Fig. 6(c) and Fig. 1a of Ref. [31]. Here, the thermal fluctuation can play an important role. Using the Ginzburg-Levanyuk criterion [42], we obtain
@@
Similar arguments apply to the square skyrmion lattice 2Q. We can “push” it to lower $T$s by increasing the $\Lambda_{\pm}$ parameter. However, thermal fluctuations are also a plausible solution, and a narrow wedge of the 2Q SkL phase at $30\lesssim T<40$ K in Fig. 6(c) can be hard to distinguish experimentally. Moreover, the topological charge of magnetic unit cells relies on the criterion $2(p+s_{\pm})>m$ derived in the continuum spin density approximation. In the lattice picture, if the difference $2(p+s_{\pm})-m\equiv\delta s\ll 1$, then for the majority of spins have $s^{z}_{i}>0$ and only a few have $s^{z}_{i}<0$. On the mean field level, it can be understood by considering the spin structure (12) with the correction (13). For $m>0$, the minima of $s^{z}_{i}$ correspond to $\mathbf{k}_{x,y}\mathbf{R}_{j}=\pi+2\pi l_{x,y}$. However, due to the incommensurate $\mathbf{k}_{x,y}$, the minimum varies among magnetic unit cells, and only part of them $\propto\delta s$ have at least one spin with $s^{z}_{i}<0$. Then, even without fluctuations, in the regions of the phase diagram close to the 2Q trivial phase, the average density of topological charge $\langle n_{\textrm{sk}}\rangle\ll 1$. In simple words, it means that “robust” skyrmions providing a tangible topological Hall effect probably do not exist in the high-$T$ part of the phase diagram, despite the corresponding phase being, in principle, stable.
VI Summary and conclusion
Using the high-temperature Landau expansion of the free energy, we analytically discuss possible noncollinear (multiple-)modulated phases of tetragonal frustrated antiferromagnets. We show that the subtle inter-phase competition can be discussed in the framework of a model that includes frustrated exchange coupling, anisotropic exchange, and/or magnetodipolar interaction, single-ion anisotropy, biquadratic exchange, and Zeeman energy. For the exchange coupling, we only assume that it has four equivalent maxima at $\pm\mathbf{k}_{x}=(\pm k,0,0)$ and $\pm\mathbf{k}_{y}=(0,\pm k,0)$ in the reciprocal space. The anisotropies are combined into a momentum-dependent biaxial term. Phase diagram topology essentially depends on the corresponding axes hierarchy.
@@ VI Summary and conclusion
Easy axes along the high-symmetry direction allow observation of a single-modulated spin-density wave 1S and helicoid 1Q, see Figs. 5(a), (b), and (c). Notably, 1Q can be stable at low $T$ only for negligible biquadratic exchange. Similarly to the in-plane easy axes case, square SkL is stable at moderate fields. Experimentally, we believe that the case of out-of-plane easy axes sheds some light on the phase diagram of GdRu2Ge2 [22]. However, the observed phase diagram in that paper is much more complicated due to a close competition between two sorts of modulation vectors, and requires a separate study.
For enhanced biquadratic exchange, we observe stabilization of SDW superpositions 2SYZ and 2SZZ at high $T$, see Figs. 6(a), (b), and (c). Notably, it can be argued that 2SYZ should emerge in GdRu2Si2 phase diagram (see Refs. [37, 31]). Then, by choosing reasonable parameters, we obtain the phase diagram shown in Fig. 6(c), which seems to reproduce all crucial observations for GdRu2Si2 [21, 31].
To conclude, by generalizing previous theoretical studies, we adapt a simple yet fruitful mean-field approach for the high-temperature part of the phase diagrams of tetragonal frustrated antiferromagnets. Interplay of complex magnetic anisotropies and biquadratic exchange leads to a variety of possible phase diagrams. Despite the square skyrmion lattice being stable even if only one of the ingredients is present, their combined effect allows for promoting it to a larger region of the phase diagram. By comparing our findings with experimental phase diagrams of relevant materials GdRu2Si2  and GdRu2Ge2, we can draw important conclusions about the corresponding microscopic parameters. In particular, for the former we predict the relative strength of the biquadratic exchange of the order of tens of percent of the corresponding Heisenberg exchange, whereas for the latter the biquadratic exchange should be negligible in comparison with the anisotropic terms, which parallels numerics of Ref. [22].
@@ Appendix A Free energies of spin structures
Here, we discuss the details of the analytical calculations and present the corresponding results. The proposed phases are mostly consist of spin-density waves (S) and screw helicoids (Q) with modulation vectors $\mathbf{k}_{x}$ and $\mathbf{k}_{y}$. This way, we can discuss all the relevant phases proposed and observed in numerical calculations [25, 27, 39, 40], as well as distinguished experimentally [21, 31]. Importantly, we consider only incommensurate spin structures, allowing us to freely choose the phases of the corresponding constituents (up to certain limitations; see discussion of the 2Q structure below). Magnetic field is oriented along $\hat{z}$ and its main effect is the homogeneous magnetization component $m$.
@@ A.2 Double-modulated superposition of spin-density waves (2S)
This spin structure can be referred to as a “magnetic bubble crystal”. It was addressed in detail in Ref. [40]. In this case,
@@ A.7 Alternative 2Q ordering (2Q)
The possibility of this spin structure emergence was predicted in Ref. [27]. In real space, it is given by
@@ Appendix B Topological properties of multiple-modulated spin structures
$\rho_{\textrm{2Q}}(\mathbf{r})=-\frac{2s^{2}\left[\sqrt{2}p\left(\cos{\mathbf{k}_{x}\mathbf{r}}+\cos{\mathbf{k}_{y}\mathbf{r}}\right)+2m\cos{\mathbf{k}_{x}\mathbf{r}}\cos{\mathbf{k}_{y}\mathbf{r}}-q\frac{6+2\cos{2\mathbf{k}_{x}\mathbf{r}}+2\cos{2\mathbf{k}_{y}\mathbf{r}}-\cos{2\left(\mathbf{k}_{x}+\mathbf{k}_{y}\right)\mathbf{r}}-\cos{2\left(\mathbf{k}_{x}-\mathbf{k}_{y}\right)\mathbf{r}}}{2\sqrt{2}}\right]}{\left[2s^{2}(\sin^{2}{\mathbf{k}_{x}\mathbf{r}}+\sin^{2}{\mathbf{k}_{y}\mathbf{r}})+\left\{2m+\sqrt{2}p(\cos{\mathbf{k}_{x}\mathbf{r}}+\cos{\mathbf{k}_{y}\mathbf{r}})-2\sqrt{2}q\cos{\mathbf{k}_{x}\mathbf{r}}\cos{\mathbf{k}_{y}\mathbf{r}}\right\}^{2}\right]^{3/2}}$
@@
We plot densities of topological charge for 1Q+1S, 2Q, and 2Q for certain reasonable parameters in Fig. 7 for a domain containing $2\times 2$ magnetic unit cells. 1Q+1S structure has zero net topological charge. However, there are peculiar stripes of topological charge due to meron-antimeron spin structure [31, 32]. Square skyrmion lattice 2Q has $n_{\textrm{sk}}=\pm 1$, whereas the alternative structure 2Q has $n_{\textrm{sk}}=\pm 2$ [25, 27].